### v4

# An Efficient Algorithm for Permutation Iteration Using a Singly Linked List

Abstract
We present a new algorithm for iterating over all permutations of a sequence. The algorithm leverages elementary $O(1)$ operations on recursive lists. As a result, no new nodes are allocated during the computation. Instead, all elements are rearranged within the original nodes of the singly linked list throughout the process. While permutations are generated in an unusual order, the transitions between consecutive permutations remain smooth. A proof of concept written in the Lisp programming language is proposed and discussed.
Keywords: permutations, singly linked list
1 Introduction
The most straightforward method for generating permutations is arguably the recursive approach. For each element $e$ in the set $A$, the permutations of the subset $A\setminus\{e\}$ are computed recursively. The element $e$ is then inserted at the beginning of each resulting sequence. Finally, the union of all computed parts is returned.
Basic implementations of this method generate all permutations in lexicographic order, but several variations can be employed. The well-known review by Robert Sedgewick enumerates these variations [2]. One of the most challenging aspects of such algorithms is selecting a new element $e$ at each step, as the recursive calls continuously reorder the list, intermingling used and unused elements.
The most commonly used data structure for the problem is an array. In this case, the new element $e_{\textrm{new}}$ is typically moved to the beginning of the sequence by swapping it with the previously selected element $e_{\textrm{prev}}$, which is a constant-time operation on arrays.
Since we focus on singly linked lists, our approach emphasizes extracting a carefully chosen element $e$ and inserting it at the beginning of the list. Furthermore, we constrain the extraction of a new element to specific locations within the list to avoid unnecessary iterations: either at the current working point or at the end of the list. As a result, the data structure is locally treated as a type of input-restricted deque in the proposed algorithm.
A surprisingly regular pattern enables elements to be popped from either the left (the current working point) or the right (the end of the list) while consistently selecting previously unused elements, despite the reordering of the list caused by recursive calls.
The skeleton of the algorithm is presented using an inefficient but expressive Python implementation. We then prove the correctness of the algorithm, primarily by demonstrating that each new element inserted at the working point was previously unused. Finally, we discuss implementation details and provide a working Lisp version of the algorithm. The Lisp programming language is chosen for its two relevant features: native support for singly linked lists and the mutability of its data structures.
While similar algorithms have been used for years, particularly within the Lisp community, they often accumulate multiple in-place modifications during the recursion, resulting in consecutive permutations that may not be closely related. In contrast, the algorithm discussed below consistently transitions between consecutive permutations by picking a single element for inserting it elsewhere.
To conclude the introduction with a practical illustration, the algorithm generates the permutations of $(0,1,2,3)$ in the following sequence:
$(0,1,2,3)$ $(0,1,3,2)$ $(0,3,1,2)$ $(0,3,2,1)$ $(0,2,3,1)$ $(0,2,1,3)$
$(2,0,1,3)$ $(2,0,3,1)$ $(2,3,0,1)$ $(2,3,1,0)$ $(2,1,3,0)$ $(2,1,0,3)$
$(3,2,1,0)$ $(3,2,0,1)$ $(3,0,2,1)$ $(3,0,1,2)$ $(3,1,0,2)$ $(3,1,2,0)$
$(1,3,2,0)$ $(1,3,0,2)$ $(1,0,3,2)$ $(1,0,2,3)$ $(1,2,0,3)$ $(1,2,3,0)$
and the distance between consecutive permutations, measured according to a relevant metric, will also be discussed.
2 Description of the algorithm
The following prototype implementation in Python emphasizes correctness and expressiveness, while disregarding efficiency and the advantages offered by more specialized data structures. It relies on the native pop and insert methods of Python lists, enabling a focus on element positions rather than low-level technical details.
1def permutations(l, i=0):
2 if i >= len(l) - 1:
3 print(l)
4 else:
5 permutations(l, i+1)
6
7 l.insert(i, l.pop(i+1))
8 permutations(l, i+1)
9
10 for _ in range(len(l)-i-3):
11 if (len(l) - i)%2 == 0: l.insert(i, l.pop())
12 else: l.insert(i, l.pop(i+1))
13 permutations(l, i+1)
14
15 if len(l) - i > 2:
16 l.insert(i, l.pop(i+1))
17 permutations(l, i+1)
When entering the permutations function, we consider the elements of the list l starting from index i, specifically the sublist l[i:] of length len(l) - i. From this point onward, let $k$ denote the number of elements in this sublist. As expected, $k$ recursive calls are made in the general case, one for each new element inserted at position i:
  • the first recursive call (at line 5) will be referred to as the “no-op” call, as it allows the termination condition to be reached without modifying the current list, thereby yielding its initial (current) permutation;
  • $k-1$ additional recursive calls are then made after inserting at position i an element that was popped either from position i+1 or from the end of the list, according to the parity of $k$ and certain positional considerations.
We refer to these recursive calls as “left” or “right”, depending on the position of the previously popped element (more precisely, each of these terms denotes the two sequential operations: the list modification and the following recursive call). The general structure of the algorithm can now be illustrated by the following triangle:
$k=2$ no-op left
$k=3$ no-op left left
$k=4$ no-op left right left
$k=5$ no-op left left left left
$k=6$ no-op left right right right left
$k=7$ no-op left left left left left left
$k=8$ no-op left right right right right right left
Although the pattern is simple and regular, it is not immediately evident that each popped element is inserted at position i for the first time, as the intermediate recursive calls significantly reorder the elements in the sublist under consideration. The next section of this paper will analyze the pattern and prove the required property.
3 Proof of correctness
To understand how the sublist is modified during a recursive call, consider the last permutation produced while iterating over all permutations of $(0,1,2,\dots,k-1)$. Although the primary purpose of the recursive call is to generate multiple permutations, it can also be viewed externally as a direct mapping of $(0,1,2,\dots,k-1)$ to a specific resulting permutation:
$k=0$ NIL
$k=1$ $(0)$
$k=2$ $(1,0)$
$k=3$ $(1,0,2)$
$k=4$ $(1,2,3,0)$
$k=5$ $(1,0,2,3,4)$
$k=6$ $(1,4,3,5,2,0)$
$k=7$ $(1,0,2,3,4,5,6)$
$k=8$ $(1,4,3,5,6,7,2,0)$
$k=9$ $(1,0,2,3,4,5,6,7,8)$
$k=10$ $(1,4,3,5,6,7,8,9,2,0)$
$k=11$ $(1,0,2,3,4,5,6,7,8,9,10)$
$k=12$ $(1,4,3,5,6,7,8,9,10,11,2,0)$
In accordance with the pattern inherent in the algorithm, two distinct cases can be easily identified based on the parity of $k$, which will now be examined in two corresponding lemmas.
Lemma 1.
Let $k$ be an odd number $(k>4)$, and assume that the last permutation of $k$ elements produced by the algorithm turns out to be the initial sequence with the first two elements swapped. When the algorithm is applied to the sequence $(0,1,2,\dots,k)$, each of the $k+1$ elements will be inserted exactly once at the beginning, and the final permutation generated by the algorithm will be $(1,4,3,5,6,7,8,\dots,k,2,0)$.
Proof.
As $k+1$ is even, the algorithm will make $k+1$ recursive calls following the pattern:
no-op, left, right, right, right, …, right, right, right, left
and according to the assumption, each recursive call will swap the second and the third elements, as the first element lies outside the scope of the recursive calls.
Thus, the initial “no-op” call will leave the sequence of $k+1$ elements as $(0,2,1,3,4,5,6,\dots,k)$, and the first “left” call will transform it into $(2,1,0,3,4,5,6,\dots,k)$. The subsequent $k-2$ “right” calls will proceed as follows:
$\begin{array}[]{l}(k,1,2,0,3,4,5,\dots,k-1)\\ (k-1,1,k,2,0,3,4,5,\dots,k-2)\\ (k-2,1,k-1,k,2,0,3,4,5,\dots,k-3)\\ \dots\\ (3,1,4,5,\dots,k,2,0)\end{array}$
Finally the “left” call will produce the expected $(1,4,3,5,6,7,8,\dots,k,2,0)$. Clearly, each of the $k+1$ elements will have been inserted exactly once at the beginning of the sequence.∎
Lemma 2.
Let’s $k$ be an even number $(k>4)$, and assume that the last permutation of $(0,1,2,\dots,k-1)$ produced by the algorithm turns out to be $(1,4,3,5,6,7,8,\dots,k-1,2,0)$. When the algorithm is applied to the sequence $(0,1,2,\dots,k)$, each of the $k+1$ elements will be inserted exactly once at the beginning, and the final permutation generated will be the initial sequence with the first two elements swapped.
Proof.
As $k+1$ is odd, the algorithm will make $k+1$ recursive calls following the pattern:
no-op, left, left, left, left, left, left, left, left…
By carefully relabeling the elements, we find that the initial “no-op” call leaves the sequence of $k+1$ elements as $(0,2,5,4,6,7,8,9,\dots,k,3,1)$.
Next, $k$ “left” recursive calls are made. However, tracking the position of each element throughout the process is not as straightforward as in the proof of Lemma 1. An simpler approach can be achieved by leveraging graph theory.
We begin by compressing the “left“ recursive call into a single permutation $T$ of the $k+1$ elements:
  • the two first elements are swapped initially;
  • the permutation of $k$ elements from the assumption is then applied to the sublist starting at the second element.
Applying $T$ to $(0,1,2,\dots,k)$ results in $(1,2,5,4,6,7,8,\dots,k,3,0)$. We now represent the list as an array and describe $T$ as a directed graph based on the following concept: $k+1$ vertices correspond to each index in the array, and $k+1$ edges illustrate how each element moves from one position to another position when $T$ is applied. The resulting graph appears as follows:
[…]
Consider an element $e$ located at position $i_{k}$; the element $e$ will jump to $i_{k-2}$, $i_{k-4}$, …, iterating over all even indices downward until it reaches $i_{4}$, and then $i_{3}$. Next, $e$ will move to $i_{k-1}$ and iterate over all odd indices downward, eventually reaching $i_{5}$. The final part of the path traverses $i_{2}$, $i_{1}$, and $i_{0}$. Observing the structure of the graph, we see that it forms a single strongly connected component, which ultimately reduces to a single directed cycle.
It is now evident that, during the $k$ iterations, all elements will pass through $i_{0}$. One additional iteration would return the element initially at $i_{0}$ back to its starting position.
To compute the last permutation of the $k+1$ elements, instead of iterating $k$ times with $T$, it is more straightforward to perform a single backward iteration. Starting from $(0,2,5,4,6,7,8,9,\dots,k,3,1)$, as mentioned above, this leads to $(1,0,2,3,4,5,\dots,k)$.
Thus, each element will be inserted at the beginning before a recursive call, and the list will return to its initial state, except for the first two elements being swapped. ∎
Theorem 3.
The algorithm described in section 2 is correct.
Proof.
The algorithm works correctly for $k$ in the range from 00 to $5$. Additionally, we verify that the last computed permutation of $(e_{0},e_{1},e_{2},e_{3},e_{4})$ is $(e_{1},e_{0},e_{2},e_{3},e_{4})$. Using Lemmas 1 and 2, we prove by induction that the algorithm selects each element of the sequence exactly once to insert it at the beginning before computing all permutations of the remaining elements. ∎
4 Transition Distance Between Consecutive Permutations
Looking at the description of the algorithm provided in Section 2 2, we can observe that no modifications occur before the first recursive call or after the last one. This leads to two resulting properties:
  • after each modification, the list is yielded without any further alterations;
  • when a recursive call returns, the list remains exactly as it was in the last yielded permutation.
The first implication is that a single modification occurs between two consecutive permutations: one element is selected from its position and inserted into another (in contrast to other similar algorithms). This property is also maintained in the odd case when cycling through all permutations, as returning to the initial permutation from the last one merely requires swapping the first two elements.
A second implication is that calculating the distance between two consecutive permutations, according to a given metric, can be achieved through purely local considerations.
The Kendall tau distance is particularly relevant in this context, as it simplifies to the length of each move in our specific case—a property that is straightforward to formalize.
Let us denote by $D_{k}$ the cumulative distance in a complete traversal of all permutations of order $k$. It is straightforward to observe that $D_{k}=kD_{k-1}+a_{k}$, where $kD_{k-1}$ represents the cumulative distance contributed by the $k$ recursive calls, and $a_{k}$ is an additional local term accounting for the cumulative distance of all modifications performed between the recursive calls.
Using the triangle that outlines the recursive calls in Section 2, we derive the following recurrence relations:
$\left\{\begin{array}[]{ll}D_{1}=0\\ D_{k}=kD_{k-1}+2+(k-3)(k-1)&\textrm{if $k$ is even}\\ D_{k}=k(k-1)D_{k-1}+k-1&\textrm{if $k$ is odd}\\ \end{array}\right.$
Solving these relations with the aid of a computer algebra system is feasible but yields cumbersome and inelegant formulas.
However, employing these formulas to calculate the average distance between two consecutive permutations for large values of $k$ yields simple expressions, as many terms vanish in the limit:
$\lim_{k\to\infty}\frac{D_{k}}{k!-1}=7\cosh(1)-4\sinh(1)-5\quad\approx 1.100759669\dots$
which appears to be a remarkably low limit for a recursive algorithm of this kind. By comparison, the average distance between two consecutive permutations in lexicographic order approaches a limit of 2.
As a final summary, each permutation is derived from the previous one by selecting an element from the current list and inserting it to the left, with an average movement of slightly more than one position.
5 Implementation technical details
We now discuss an implementation of the algorithm using singly linked lists. Such a data structure is typically not well-suited for extracting elements repeatedly from both the beginning and the end of a sublist. However, in this very specific case, a natural and elegant solution to this issue will emerge. We highlight that making this extraction an atomic operation is one of the key features of our algorithm.
We also require the data structure to be mutable, enabling a highly efficient implementation of our algorithm by moving and reusing nodes without creating new ones (eliminating the need for “cons” operations). This requirement led us to choose Lisp for the proof of concept; however, nothing in the code depends on Lisp-specific features, and it can easily be adapted to other programming languages. This second requirement is, of course, incompatible with working in a purely functional style, as described in Okasaki’s renowned book [1].
Iterating over the consecutive sublists is very natural with such a data structure (by repeatedly taking the “cdr” of the list), and swapping the two initial elements at a given location is straightforward.
We also note that extracting the last element of the sublist is never done before a recursive call, which leads to the following idea:
  • the descending phase of the recursive call is intended to compute new partial permutations;
  • the ascending phase of the recursive call is intended to pass information needed by the next recursive call (returning a handle to the end of the list).
The figure below illustrates the “current” frame near the middle of the recursive call stack. An undefined element is added at the beginning of the list to ensure consistency across all calls. At this level, the sublist operates as an input-restricted deque:
$e_{0}$$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$$e_{6}$$e_{7}$$e_{8}$$e_{9}$current framehandle to the beginning of the listhandle to the current “head” nodecurrent “head” nodehandle to theend of the list(return value)input-restricted deque
Full advantage is taken here of the observation made in Section 4: only a single alteration is made to the list between two consecutive permutations. Consequently, the handle to the last element is used once, becoming immediately obsolete afterward, and no additional handle is needed until the next permutation is generated.
The following code provides a highly efficient implementation of the algorithm:
1(defun permutations (li f)
2 (labels ((aux-odd (p func n handle)
3 (loop with s = (1- n)
4 initially (aux-even (cdr p) func s p handle)
5 repeat n
6 do (progn (rotatef (car p) (cadr p))
7 (aux-even (cdr p) func s p handle))
8 finally (return (progn
9 (rotatef (car p) (cadr p))
10 (aux-even (cdr p) func s p handle)))))
11 (aux-even (p func n prev handle)
12 (if (cddr p)
13 (loop with s = (1- n)
14 initially (progn (aux-odd (cdr p) func s handle)
15 (rotatef (car p) (cadr p)))
16 repeat s
17 do (let ((l (aux-odd (cddr prev) func s handle)))
18 (rotatef (cddr l) (cdr prev) (cdr l)))
19 finally (return (progn
20 (setf p (cdr prev))
21 (aux-odd (cdr p) func s handle)
22 (rotatef (car p) (cadr p))
23 (aux-odd (cdr p) func s handle))))
24 (progn (funcall func (cdr handle))
25 (rotatef (car p) (cadr p))
26 (funcall func (cdr handle)) p))))
27 (let ((q (cons NIL li)) (n (length li)))
28 (if (> n 1)
29 (if (oddp n) (aux-odd li f (- n 2) q)
30 (aux-even li f (- n 2) q q))
31 (funcall f li)) NIL)))
The function performs a single “cons” operation (at line 27) to create a handle pointing to the actual beginning of the list, as the initial node may be repositioned during the process. The two auxiliary mutually recursive functions are subsequently invoked with either four or five arguments, depending on the parity of the current length:
  • the considered list, p;
  • the callback function func, to be applied to each generated permutation;
  • the length of the sublist decreased by 2 (in order to match the size of the loops without changing its parity);
  • a handle to the previous node (before p) called prev, which always exists, since even the first useful element of the list has a preceding (NIL) element from the initial “cons” at line 27;
  • a handle (also called handle) to the beginning of the extended list.
The recursive functions return a handle to the penultimate node of the list, which allows extracting the last node to insert it before the current head node.
We chose $k=2$ for the halting case, as this specific case required slightly different handling compared to other even cases. Keeping the cases $k=0$ and $k=1$ outside the recursive process was straightforward.
The most important point to note here is that, apart from the recursive calls, the only operations performed by the function are elementary assignments using the setf or rotatef macros.
Benchmarking suggests that the previous code achieves performance comparable to classical array-based recursive algorithms while preserving the list data type throughout.
6 Conclusion
In this paper, we introduced a novel algorithm for generating permutations using singly linked lists, emphasizing efficiency and elegance. Unlike conventional methods that rely on arrays and repeated node allocations, our approach leverages the mutability of linked lists to rearrange nodes dynamically without additional memory overhead. By utilizing a recursive structure and distinguishing between the descending and ascending phases of recursive calls, the algorithm efficiently handles both permutation generation and the propagation of structural information about the list.
This work illustrates the potential for alternative data structures, such as singly linked lists, to solve traditionally array-dominated problems in a computationally efficient and conceptually elegant way. Future work may explore extending this approach to other combinatorial problems.
Declarations
The author has no competing interests to declare that are relevant to the content of this article.